OpenStudy (anonymous):

lim of sin x / x x -> 0

OpenStudy (anonymous):

l'Hopital's rule, yo.\[\lim_{x\to0}\frac{\sin x}{x}=\lim_{x\to0}\frac{\cos x}{1}\]

OpenStudy (anonymous):

how did u do that?

OpenStudy (anonymous):

you are right . limit is = 1 but how did u do that ?

OpenStudy (anonymous):

Let's say we wish to compute\[\lim_{x\to a}\frac{f(x)}{g(x)}\]and we know\[\left\{\lim_{x\to a}f(x)=\lim_{x\to a}g(x)=\infty,0\right\}\forall a\in\mathbb R\]then\[\lim_{x\to a}\frac{f(x)}{g(x)}=\lim_{x\to a}\frac{f'(x)}{g'(x)}\]